OpenStudy (angel_kitty12):

Please help with probability. I'm really bad at it and I really need to study this but I can't work around the problem. Help please. Thank you so much. If a safety switch has a 1 in 10 chance of failing, how many switches would a company want to install in order to have only a 1 in one million chance of them all failing at the same time? Show your reasoning.

OpenStudy (jack1):

P(1) = 1/10 P(1 AND 2 AND 3 AND....n) = 1/1000000 n = ?

OpenStudy (angel_kitty12):

Would it be 1/10•1/10•1/10 ?

OpenStudy (jack1):

if it helps P(1 AND 2 AND 3 AND....n) = 1/1000000 P (1) = 1/10 P (2) = 1/10 P (3) = 1/10 so P(1 AND 2 AND 3) = P(1) x P(2) x P(3) = 1/10 x 1/10 x 1/10 = (1/10)^3 = 1/1000 note... 1/1000000 = (1/10)^6

OpenStudy (angel_kitty12):

No 1/10•1/10•1/10•1/10•1/10•1/10 6?

OpenStudy (angel_kitty12):

Okay I understand hat so would the answer be 6?